# Optimizing Fine-Grained Parallelism Through Dynamic Load Balancing on Multi-Socket Many-Core Systems

### v1

@@ IX Related Work
Dynamic Load balancing: Several papers have proposed load balancing mechanisms [35, 36]. Quintin et al. proposed hierarchical work stealing to exploit data locality to achieve speedup over classical work stealing algorithms [37]. Shiina et al. addressed the issue of data locality by making scheduling deterministic [38]. In order to balance load efficiently, these all relied on synchronisation mechanisms. In our work, we explore lock-less techniques to achieve comparable dynamic load balancing mechanisms.
@@ X Discussion and Future Work
GNU OpenMP, although part of the mainstream compiler infrastructure GCC, is unable to harness modern high-performance CPUs with ever-increasing number of cores, due to its conservative design of using excessive locks and atomic operations. Our work transforms GNU OpenMP into a high-performance parallel computing library by integrating the novel lock-less concurrent data structure XQueue and an efficient distributed tree barrier to achieve up to 1522.8$\times$ performance improvement compared to the original GNU OpenMP. We further improve performance via use of dynamic load balancing strategies, demonstrating that with optimal settings, lock-less work stealing can achieve 4$\times$ more performance improvement compared to use of a static load balancer.

### v2

@@ IX Related Work
Dynamic Load balancing: Several papers have proposed load balancing mechanisms [35, 36]. Quintin et al. proposed hierarchical work stealing to exploit data locality to achieve speedup over classical work stealing algorithms [37]. Shiina et al. addressed the issue of data locality by making scheduling deterministic [38]. In order to balance load efficiently, these all relied on synchronization mechanisms. In our work, we explore lock-less techniques to achieve comparable dynamic load balancing mechanisms.
@@ X Discussion and Future Work
GNU OpenMP, although part of the mainstream compiler infrastructure GCC, is unable to harness modern high-performance CPUs with many cores, due to its use of locks and atomic operations. Our work improves GNU OpenMP performance by integrating a novel lock-less concurrent data structure, XQueue, and an efficient distributed tree barrier to achieve up to 1522.8$\times$ improvement compared to the original GNU OpenMP. We further improve performance through the use of dynamic load balancing strategies, demonstrating that with optimal settings, lock-less dynamic load balancing can achieve 4$\times$ more performance improvement compared to using static load balancing.
@@
Acknowledgments
This work is supported in part by the NSF CNS-1730689 CRI award (Mystic Testbed [39]), as well as the NSF OAC-2107283 and OAC-2107548 awards.